9. Use the Laplace transform to solve the system dx -xty dt dy dt x(0) = 0, y(0) = 1 = 2x
Answer #1

b Given system is. arty = coin o du at dy at Y10) 22 we can write. st u ih -xty 220 = ) taking the Laplace transform both equmultiply 0 by Sist) x1y) -gy) =0 3 add 2 4 3 (515+1) – 2 ) xls) X (5) = s² +5-2 Xit) = Lt { taking inverse Laplace transform크 () 2 + 4-2 taking inverse Laplace transform. 개 = 네 커 2 +스 ) 2. } ) 1242)} 를녜 급 }같네 네 3(S) 2 2+ y lt) 2 르 et + 3 CS Scanned